# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v6

@@ IV-A1 $\mathbf{C}_{0}$-Vanilla Prompting
TABLE II: A summary of representative methods used in the ICL paradigm for LLM-based text-to-SQL is provided. The complete table, along with a detailed categorization of existing well-designed methods denoted as $\mathbf{C}_{1:4}$, is available in Tab. III. *Note that SimpleDDL is a prompting style that provides only table and column names [120], rather than a tailored method.
@@
$\mathbf{C}_{0}$-Vanilla Prompting SimpleDDL [120]* SQLCoder
$\mathbf{C}_{1}$-Decomposition DIN-SQL [7] GPT-4
$\mathbf{C}_{2}$-Prompt Optimization DAIL-SQL [8] GPT-4
$\mathbf{C}_{3}$-Reasoning Enhancement ACT-SQL [68] GPT-4
$\mathbf{C}_{4}$-Execution Refinement LEVER [79] Codex
@@
TABLE III: Well-designed methods in the ICL paradigm for LLM-based text-to-SQL are ordered by release time. The methods are grouped into four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement. Methods that fall into multiple categories will be introduced accordingly. The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [11], BIRD [1], Spider-DK [36] Spider-SYN [37], Spider-Realistic [38], and Spider 2.0 [42]
@@
. We use the reference ID in the table due to column space limitations. *There are multiple applied LLMs in the corresponding method; the model with representative performance is presented. CoT method is introduced in multiple venues: NeurIPS’23 [1], EMNLP’23 [47], VLDB’24 [8], arXiv’24 [30]. Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue MBR-Exec [78] Codex [11] EX Apr-2022 EMNLP’22 Coder-Reviewer [46] Codex [11] EX Nov-2022 ICML’23 LEVER [79] Codex [11] EX Feb-2023 ICML’23 Self-Debugging [80] StarCoder* [11] EX Apr-2023 ICLR’24 DESEM+P [63] ChatGPT [11, 37] EX Apr-2023 PRICAI’23 DIN-SQL [7] GPT-4* [1, 11, 42] EX, EM, VES Apr-2023 NeurIPS’23 CoT [47, 1, 8, 30] GPT-4 [11, 38, 1] EX, VES May-2023 Multiple Venues StructGPT [64] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23 SD+SA+Voting [65] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23 Findings QDecomp [47] Codex [11, 38] EX May-2023 EMNLP’23 Least-to-Most [47] Codex [11] EX May-2023 EMNLP’23 SQL-PaLM [74] PaLM-2 [11] EX May-2023 TMLR’24 RAG+SP&DRC [66] ChatGPT [11] EX Jul-2023 ICONIP’23 C3 [28] ChatGPT [11] EX Jul-2023 arXiv’23 DAIL-SQL [8] GPT-4* [11, 1, 38, 42] EX, EM, VES Aug-2023 VLDB’24 ODIS [67] Codex* [11] EX Oct-2023 EMNLP’23 Findings ACT-SQL [68] GPT-4* [11, 37] EX, EM Oct-2023 EMNLP’23 Findings MAC-SQL [48] GPT-4* [11, 1] EX, EM, VES Dec-2023 COLING’25 DEA-SQL [49] GPT-4 [11] EX Feb-2024 ACL’24 Findings FUSED [69] ChatGPT* [11] EX Feb-2024 EMNLP’24 Findings DELLM [70] GPT-4* [11, 1] EX, VES Feb-2024 ACL’24 Findings SGU-SQL [30] GPT-4* [11, 1] EX, EM Feb-2024 ICML’25 PoT [75] GPT-4* [11, 1] EX Feb-2024 arXiv’24 SQL-CRAFT [75] GPT-4* [11, 1] EX Feb-2024 arXiv’24 FUXI [76] GPT-4* [1] EX Feb-2024 EMNLP’24 MetaSQL [50] GPT-4* [11] EX, EM Feb-2024 ICDE’24 $\textit{{R}}^{3}$ [51] GPT-4* [11, 1] EX Feb-2024 ACL’25 Workshop PET-SQL [52] GPT-4 [11] EX Mar-2024 arXiv’24 PURPLE [53] GPT-4* [11, 37, 38] EX, EM Mar-2024 ICDE’24 Dubo-SQL [71] ChatGPT* [1] EX Apr-2024 arXiv’24 TA-SQL [54] GPT-4* [11, 1, 36, 38] EX Apr-2024 ACL’24 Findings MCS-SQL [55] GPT-4 [11, 1] EX, VES May-2024 COLING’25 CHESS [56] GPT-4o* [11, 1, 42] EX, VES May-2024 arXiv’24 SuperSQL [57] GPT-4* [11, 1] EX, EM, VES Jun-2024 VLDB’24 Distillery [58] GPT-4o-Mini* [1] EX Aug-2024 NeurIP’24 Workshop E-SQL [81] GPT-4o* [11, 1] EX, VES Sep-2024 arXiv’24 CHASE-SQL [77] Gemini-1.5* [11, 1] EX, VES Oct-2024 ICLR’25 RSL-SQL [59] DeepSeek-V3* [11, 1] EX, VES Nov-2024 arXiv’24 Spider-Agent [42] o1-preview* [42] EX Nov-2024 ICLR’25 ICRL [72] ChatGPT [11] EX Jan-2025 COLING’25 Gen-SQL [60] LLaMA-3* [11, 1] EX, EM, VES Jan-2025 COLING’25 SAFE-SQL [73] GPT-4o-Mini* [11] EX, EM Feb-2025 arXiv’25 ReFoRCE [61] o1-preview* [42] EX Feb-2025 ICLR’25 Workshop LinkAlign [62] DeepSeek-R1* [1, 42] EX, EM Mar-2025 arXiv’25
@@
1. Zero-shot studies: The zero-shot-based ICL strategies primarily focus on studying the influence of prompt representation styles and the baseline performance of various LLMs for text-to-SQL [6, 25, 120]. As an empirical evaluation, [6] evaluates the baseline text-to-SQL capabilities of various early-developed LLMs [123, 106, 124] and their performance across different prompting styles. The experimental results indicate that prompt design is critical for performance. Through error analysis, this study [6] further suggests that including more database content can harm the overall accuracy of text-to-SQL. Since the emergence of ChatGPT [106] with its impressive capabilities in conversational scenarios and code generation [125], [25] evaluates its performance on Spider [11] and its variant [37, 36, 38] datasets. Under zero-shot settings, the results demonstrate that ChatGPT has promising text-to-SQL performance compared to state-of-the-art PLM-based systems. Aside from the proprietary models discussed above, similar evaluations indicate that open-source LLMs can also perform zero-shot text-to-SQL tasks [114, 83, 120], particularly for the code-specific models [80, 120].
@@
Primary keys and foreign keys convey continuous knowledge across different tables, making them an effective prompting style in zero-shot scenarios. [68] studied their impact by incorporating these keys into various prompt representation styles with different database content to analyze performance outcomes. A benchmark evaluation [8] also assessed the influence of foreign keys using five distinct prompt representation styles. Each style can be viewed as a permutation and combination of instruction, rule implication, and foreign key elements. In addition to using foreign keys, this study explores zero-shot prompting combined with the rule implication “no explanation” to produce concise outputs. With the help of annotated external knowledge from human experts, the evaluation in the BIRD benchmark [1] significantly improves by incorporating the provided annotated oracle knowledge.
@@
In terms of zero-shot prompting optimization, [120] highlights the challenge of designing effective prompt templates for LLMs. Previous simple prompt constructions often lack structural uniformity, complicating the identification of specific elements in a prompt template that influence LLM performance. This challenge is addressed by investigating a more unified set of prompt templates that include various prefixes, infixes, and postfixes. To ensure fair comparability, [108] explores effective prompt construction for LLMs by studying different styles of prompt representation and highlights their findings on the crucial role of table structure, content, prompt length, and optimal representations in effective prompting.
@@
2. Few-shot studies: The study of few-shot prompting strategies focuses on the number and selection of few-shot instances. Empirical studies have evaluated few-shot prompting for text-to-SQL across multiple datasets using various LLMs [7, 30], demonstrating solid improvements compared to zero-shot prompting. [1] provides a detailed 1-shot example to trigger the text-to-SQL model for generating accurate SQL. [75] explores the effect of the number of few-shot examples. [65] focuses on sampling strategies by examining the similarity and diversity between different demonstrations, setting random sampling as the baseline, and evaluating various strategies and their combinations for comparison. Furthermore, beyond similarity-based selection, [8] evaluates masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples.
@@
A study on difficulty-level sample selection [47] compared the performance of few-shot Codex [124], using random selection and difficulty-based selection for few-shot instances on difficulty-categorized datasets [11, 38]. Three difficulty-based selection strategies are devised based on samples selected by difficulty level. [68] utilizes a hybrid strategy for selecting samples, combining static examples and similarity-based dynamic examples for few-shot prompting. In their settings, they also evaluate the impact of different input schema styles and various numbers of static and dynamic exemplars.
@@
The impact of cross-domain examples is also being studied [67]. When incorporating in- and out-of-domain examples in different numbers, the in-domain demonstrations outperform zero-shot and out-of-domain examples, with performance improving as the number of examples increases. To explore the detailed construction of prompts, [74] compares concise and verbose prompt design strategies. The former style separates the schema, column names, primary keys, and foreign keys, while the latter organizes them as natural language descriptions.
@@ IV-A2 $\mathbf{C}_{1}$-Decomposition
1. Multi-module collaboration: This strategy breaks the task into specialized stages (e.g., schema linking, SQL generation, and SQL refinement) handled by distinct modules or agents. C3 [28] employs a three-module pipeline involving schema linking via clear prompting, calibration bias prompting for SQL guidance, and consistency-based voting for final selection. Similarly, MAC-SQL [48] and CHESS [56] use multi-agent frameworks where dedicated agents handle schema selection, decomposition, and refinement. Workflow-centric methods, like Distillery [58] and DEA-SQL [49], decompose the task into retrieval, generation, and correction stages. Iterative refinement and structural decomposition are also prevalent. DIN-SQL [7] integrates a four-stage pipeline, including schema linking, classification, SQL generation, and self-correction. Coder-Reviewer [46] separates generation and validation into distinct models. PET-SQL [52] refines outputs through multi-stage prompting, while MCS-SQL [55] uses execution-guided selection as the final module after multiple SQL generation. With the increasing complexity of recently proposed benchmarks, agentic frameworks like ReFoRCE [61], LinkAlign [62] and Spider-Agent [42] leveraged well-designed step-by-step action plans for LLMs to systematically reason toward the final SQL.
2. Question decomposition: Complexity is reduced by splitting user questions into intermediate sub-questions via question decomposition. QDecomp [47] uses Least-to-Most prompting, breaking complex questions into step-by-step sub-questions to facilitate SQL generation. SGU-SQL [30] employs graph-based structure linking and meta-operators for decomposing grammar-aware sub-questions. Metadata-driven approaches, such as MetaSQL [50] and TA-SQL [54], further decompose questions using metadata conditioning and task-aligned schema linking. Meanwhile, frameworks like DIN-SQL [7] and MAC-SQL [48] use question decomposition as auxiliary modules but primarily focus on module-level collaboration.

### v7

@@ IV-A1 $\mathbf{C}_{0}$-Vanilla Prompting
TABLE II: A summary of representative methods used in the ICL paradigm for LLM-based text-to-SQL is provided. The complete table, along with a detailed categorization of existing well-designed methods denoted as $\mathbf{C}_{1:4}$, is available in Tab. III. *Note that SimpleDDL is a prompting style that provides only table and column names [zhang2024benchmarking], rather than a tailored method.
@@
$\mathbf{C}_{0}$-Vanilla Prompting SimpleDDL [zhang2024benchmarking]* SQLCoder
$\mathbf{C}_{1}$-Decomposition DIN-SQL [pourreza2023dinsql] GPT-4
$\mathbf{C}_{2}$-Prompt Optimization DAIL-SQL [gao2023dailsql] GPT-4
$\mathbf{C}_{3}$-Reasoning Enhancement ACT-SQL [zhang2023actsql] GPT-4
$\mathbf{C}_{4}$-Execution Refinement LEVER [ni2023lever] Codex
@@
TABLE III: Well-designed methods in the ICL paradigm for LLM-based text-to-SQL are ordered by release time. The methods are grouped into four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement. Methods that fall into multiple categories will be introduced accordingly. The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [yu2018Spider], BIRD [li2023BIRD], Spider-DK [gan2021Spider-DK] Spider-SYN [gan2021Spider-SYN], Spider-Realistic [deng2021Spider-Realistic], and Spider 2.0 [lei2025spider2]
@@
. We use the reference ID in the table due to column space limitations. *There are multiple applied LLMs in the corresponding method; the model with representative performance is presented. CoT method is introduced in multiple venues: NeurIPS’23 [li2023BIRD], EMNLP’23 [tai2023exploring], VLDB’24 [gao2023dailsql], arXiv’24 [zhang2024sgusql]. Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue MBR-Exec [shi2022natural] Codex [yu2018Spider] EX Apr-2022 EMNLP’22 Coder-Reviewer [zhang2022coder] Codex [yu2018Spider] EX Nov-2022 ICML’23 LEVER [ni2023lever] Codex [yu2018Spider] EX Feb-2023 ICML’23 Self-Debugging [chen2024selfdebugging] StarCoder* [yu2018Spider] EX Apr-2023 ICLR’24 DESEM+P [guo2023desem] ChatGPT [yu2018Spider, gan2021Spider-SYN] EX Apr-2023 PRICAI’23 DIN-SQL [pourreza2023dinsql] GPT-4* [li2023BIRD, yu2018Spider, lei2025spider2] EX, EM, VES Apr-2023 NeurIPS’23 CoT [tai2023exploring, li2023BIRD, gao2023dailsql, zhang2024sgusql] GPT-4 [yu2018Spider, deng2021Spider-Realistic, li2023BIRD] EX, VES May-2023 Multiple Venues StructGPT [jiang2023structgpt] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23 SD+SA+Voting [nan2023enhancing] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23 Findings QDecomp [tai2023exploring] Codex [yu2018Spider, deng2021Spider-Realistic] EX May-2023 EMNLP’23 Least-to-Most [tai2023exploring] Codex [yu2018Spider] EX May-2023 EMNLP’23 SQL-PaLM [sun2023sqlpalm] PaLM-2 [yu2018Spider] EX May-2023 TMLR’24 RAG+SP&DRC [guo2023ragsql] ChatGPT [yu2018Spider] EX Jul-2023 ICONIP’23 C3 [dong2023c3] ChatGPT [yu2018Spider] EX Jul-2023 arXiv’23 DAIL-SQL [gao2023dailsql] GPT-4* [yu2018Spider, li2023BIRD, deng2021Spider-Realistic, lei2025spider2] EX, EM, VES Aug-2023 VLDB’24 ODIS [chang2023selective] Codex* [yu2018Spider] EX Oct-2023 EMNLP’23 Findings ACT-SQL [zhang2023actsql] GPT-4* [yu2018Spider, gan2021Spider-SYN] EX, EM Oct-2023 EMNLP’23 Findings MAC-SQL [wang2024macsql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Dec-2023 COLING’25 DEA-SQL [xie2024deasql] GPT-4 [yu2018Spider] EX Feb-2024 ACL’24 Findings FUSED [wang2024fused] ChatGPT* [yu2018Spider] EX Feb-2024 EMNLP’24 Findings DELLM [hong2024knowledgetosql] GPT-4* [yu2018Spider, li2023BIRD] EX, VES Feb-2024 ACL’24 Findings SGU-SQL [zhang2024sgusql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM Feb-2024 ICML’25 PoT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24 SQL-CRAFT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24 FUXI [gu2024fuxi] GPT-4* [li2023BIRD] EX Feb-2024 EMNLP’24 MetaSQL [fan2024metasql] GPT-4* [yu2018Spider] EX, EM Feb-2024 ICDE’24 $\textit{{R}}^{3}$ [xia2024r3] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 ACL’25 Workshop PET-SQL [li2024petsql] GPT-4 [yu2018Spider] EX Mar-2024 arXiv’24 PURPLE [ren2024purple] GPT-4* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX, EM Mar-2024 ICDE’24 Dubo-SQL [thorpe2024dubo] ChatGPT* [li2023BIRD] EX Apr-2024 arXiv’24 TA-SQL [qu2024generation] GPT-4* [yu2018Spider, li2023BIRD, gan2021Spider-DK, deng2021Spider-Realistic] EX Apr-2024 ACL’24 Findings MCS-SQL [lee2025mcs] GPT-4 [yu2018Spider, li2023BIRD] EX, VES May-2024 COLING’25 CHESS [talaei2024chess] GPT-4o* [yu2018Spider, li2023BIRD, lei2025spider2] EX, VES May-2024 arXiv’24 SuperSQL [li2024dawn] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Jun-2024 VLDB’24 Distillery [maamari2024the] GPT-4o-Mini* [li2023BIRD] EX Aug-2024 NeurIP’24 Workshop E-SQL [caferouglu2024esql] GPT-4o* [yu2018Spider, li2023BIRD] EX, VES Sep-2024 arXiv’24 CHASE-SQL [pourreza2025chasesql] Gemini-1.5* [yu2018Spider, li2023BIRD] EX, VES Oct-2024 ICLR’25 RSL-SQL [cao2024rsl] DeepSeek-V3* [yu2018Spider, li2023BIRD] EX, VES Nov-2024 arXiv’24 Spider-Agent [lei2025spider2] o1-preview* [lei2025spider2] EX Nov-2024 ICLR’25 ICRL [toteja2025context] ChatGPT [yu2018Spider] EX Jan-2025 COLING’25 Gen-SQL [shi2025gen] LLaMA-3* [yu2018Spider, li2023BIRD] EX, EM, VES Jan-2025 COLING’25 SAFE-SQL [lee2025safe] GPT-4o-Mini* [yu2018Spider] EX, EM Feb-2025 arXiv’25 ReFoRCE [deng2025reforce] o1-preview* [lei2025spider2] EX Feb-2025 ICLR’25 Workshop LinkAlign [wang2025linkalign] DeepSeek-R1* [li2023BIRD, lei2025spider2] EX, EM Mar-2025 arXiv’25
@@
1. Zero-shot studies: The zero-shot-based ICL strategies primarily focus on studying the influence of prompt representation styles and the baseline performance of various LLMs for text-to-SQL [rajkumar2022evaluating, liu2023comprehensive, zhang2024benchmarking]. As an empirical evaluation, [rajkumar2022evaluating] evaluates the baseline text-to-SQL capabilities of various early-developed LLMs [raffel2020exploring, brown2020gpt3, chen2021evaluating] and their performance across different prompting styles. The experimental results indicate that prompt design is critical for performance. Through error analysis, this study [rajkumar2022evaluating] further suggests that including more database content can harm the overall accuracy of text-to-SQL. Since the emergence of ChatGPT [brown2020gpt3] with its impressive capabilities in conversational scenarios and code generation [ray2023chatgpt], [liu2023comprehensive] evaluates its performance on Spider [yu2018Spider] and its variant [gan2021Spider-SYN, gan2021Spider-DK, deng2021Spider-Realistic] datasets. Under zero-shot settings, the results demonstrate that ChatGPT has promising text-to-SQL performance compared to state-of-the-art PLM-based systems. Aside from the proprietary models discussed above, similar evaluations indicate that open-source LLMs can also perform zero-shot text-to-SQL tasks [xue2024dbgpt, xu2024symbolllm, zhang2024benchmarking], particularly for the code-specific models [chen2024selfdebugging, zhang2024benchmarking].
@@
Primary keys and foreign keys convey continuous knowledge across different tables, making them an effective prompting style in zero-shot scenarios. [zhang2023actsql] studied their impact by incorporating these keys into various prompt representation styles with different database content to analyze performance outcomes. A benchmark evaluation [gao2023dailsql] also assessed the influence of foreign keys using five distinct prompt representation styles. Each style can be viewed as a permutation and combination of instruction, rule implication, and foreign key elements. In addition to using foreign keys, this study explores zero-shot prompting combined with the rule implication “no explanation” to produce concise outputs. With the help of annotated external knowledge from human experts, the evaluation in the BIRD benchmark [li2023BIRD] significantly improves by incorporating the provided annotated oracle knowledge.
@@
In terms of zero-shot prompting optimization, [zhang2024benchmarking] highlights the challenge of designing effective prompt templates for LLMs. Previous simple prompt constructions often lack structural uniformity, complicating the identification of specific elements in a prompt template that influence LLM performance. This challenge is addressed by investigating a more unified set of prompt templates that include various prefixes, infixes, and postfixes. To ensure fair comparability, [chang2023how] explores effective prompt construction for LLMs by studying different styles of prompt representation and highlights their findings on the crucial role of table structure, content, prompt length, and optimal representations in effective prompting.
@@
2. Few-shot studies: The study of few-shot prompting strategies focuses on the number and selection of few-shot instances. Empirical studies have evaluated few-shot prompting for text-to-SQL across multiple datasets using various LLMs [pourreza2023dinsql, zhang2024sgusql], demonstrating solid improvements compared to zero-shot prompting. [li2023BIRD] provides a detailed 1-shot example to trigger the text-to-SQL model for generating accurate SQL. [xia2024sqlcraft] explores the effect of the number of few-shot examples. [nan2023enhancing] focuses on sampling strategies by examining the similarity and diversity between different demonstrations, setting random sampling as the baseline, and evaluating various strategies and their combinations for comparison. Furthermore, beyond similarity-based selection, [gao2023dailsql] evaluates masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples.
@@
A study on difficulty-level sample selection [tai2023exploring] compared the performance of few-shot Codex [chen2021evaluating], using random selection and difficulty-based selection for few-shot instances on difficulty-categorized datasets [yu2018Spider, deng2021Spider-Realistic]. Three difficulty-based selection strategies are devised based on samples selected by difficulty level. [zhang2023actsql] utilizes a hybrid strategy for selecting samples, combining static examples and similarity-based dynamic examples for few-shot prompting. In their settings, they also evaluate the impact of different input schema styles and various numbers of static and dynamic exemplars.
@@
The impact of cross-domain examples is also being studied [chang2023selective]. When incorporating in- and out-of-domain examples in different numbers, the in-domain demonstrations outperform zero-shot and out-of-domain examples, with performance improving as the number of examples increases. To explore the detailed construction of prompts, [sun2023sqlpalm] compares concise and verbose prompt design strategies. The former style separates the schema, column names, primary keys, and foreign keys, while the latter organizes them as natural language descriptions.
@@ IV-A2 $\mathbf{C}_{1}$-Decomposition
1. Multi-module collaboration: This strategy breaks the task into specialized stages (e.g., schema linking, SQL generation, and SQL refinement) handled by distinct modules or agents. C3 [dong2023c3] employs a three-module pipeline involving schema linking via clear prompting, calibration bias prompting for SQL guidance, and consistency-based voting for final selection. Similarly, MAC-SQL [wang2024macsql] and CHESS [talaei2024chess] use multi-agent frameworks where dedicated agents handle schema selection, decomposition, and refinement. Workflow-centric methods, like Distillery [maamari2024the] and DEA-SQL [xie2024deasql], decompose the task into retrieval, generation, and correction stages. Iterative refinement and structural decomposition are also prevalent. DIN-SQL [pourreza2023dinsql] integrates a four-stage pipeline, including schema linking, classification, SQL generation, and self-correction. Coder-Reviewer [zhang2022coder] separates generation and validation into distinct models. PET-SQL [li2024petsql] refines outputs through multi-stage prompting, while MCS-SQL [lee2025mcs] uses execution-guided selection as the final module after multiple SQL generation. With the increasing complexity of recently proposed benchmarks, agentic frameworks like ReFoRCE [deng2025reforce], LinkAlign [wang2025linkalign] and Spider-Agent [lei2025spider2] leveraged well-designed step-by-step action plans for LLMs to systematically reason toward the final SQL.
2. Question decomposition: Complexity is reduced by splitting user questions into intermediate sub-questions via question decomposition. QDecomp [tai2023exploring] uses Least-to-Most prompting, breaking complex questions into step-by-step sub-questions to facilitate SQL generation. SGU-SQL [zhang2024sgusql] employs graph-based structure linking and meta-operators for decomposing grammar-aware sub-questions. Metadata-driven approaches, such as MetaSQL [fan2024metasql] and TA-SQL [qu2024generation], further decompose questions using metadata conditioning and task-aligned schema linking. Meanwhile, frameworks like DIN-SQL [pourreza2023dinsql] and MAC-SQL [wang2024macsql] use question decomposition as auxiliary modules but primarily focus on module-level collaboration.